### v2

Abstract
Let $D(G)=(d_{ij})_{n\times n}$ be the distance matrix of a graph $G$ of order $n$. The largest eigenvalue of $D(G)$, denoted by $\mu(G)$, is called the distance spectral radius of $G$. A fractional matching of $G$ is a function $f:E(G)\to[0,1]$ such that for any $v_{i}\in V(G)$, $\sum_{e\in E_{G}(v_{i})}f(e)\leq 1$, where $E_{G}(v_{i})=\{e:e\in E(G)\ \textrm{and}\ e\ \textrm{is incident with}\ v_{i}\}$. Let $\alpha_{f}(G)$ denote the fractional matching number of $G$ which is defined as $\alpha_{f}(G)=\max\{\sum_{e\in E(G)}f(e):f\ \textrm{is a fractional matching of}\ G\}$. In this paper, we establish a sharp upper bound for the distance spectral radius to guarantee $\alpha_{f}(G)>\frac{n-k}{2}$ in graph $G$ with given minimum degree $\delta$.
Let $\{G_{1},G_{2},G_{3},\cdots\}$ be a set of graphs, an $\{G_{1},G_{2},G_{3},\cdots\}$-factor of a graph $G$ is a spanning subgraph of $G$ such that each component of which is isomorphic to one of $\{G_{1},G_{2},G_{3},\cdots\}$. In this paper, we give a sharp upper bound on distance spectral radius of graph $G$ with given minimum degree $\delta$ to ensure that $G$ has a $\{K_{2},C_{k}\}$-factor, where $k\geq 3$ is an integer. In addition, we also obtain a sharp upper bound on distance spectral radius for the existence of a $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor with $k\geq 2$ in a graph $G$ with given minimum degree $\delta$.
Key Words:  Fractional matching, Distance spectral radius, Factor, Minimum degree. AMS Subject Classification (2000):  05C50,15A18
1 Introduction
Throughout this paper all graphs considered are simple, connected and undirected. A graph $G$ is denoted by $G=(V(G),E(G))$, where $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$ is the vertex set and $E(G)$ is the edge set. The order of $G$ is $|V(G)|=n$ and its size is $|E(G)|=m$. The set of neighbors of the vertex $v_{i}$ denoted by $N_{G}(v_{i})$, is the set of vertices adjacent to $v_{i}$. The degree of the vertex $v_{i}$ in $G$ denoted by $d(v_{i})$, is the number of vertices of $G$ adjacent to $v_{i}$, i.e. $d(v_{i})=|N_{G}(v_{i})|$. Let $\Delta$ and $\delta$ denote the maximum degree and the minimum degree of $G$, respectively. Let $C_{n}$, $K_{n}$ and $K_{1,n}$ denote the cycle of order $n$, the complete graph of order $n$ and the star with $n$ vertices, respectively. For two vertex disjoint graphs $G_{1}$ and $G_{2}$, we use $G_{1}+G_{2}$ denote the disjoint union of $G_{1}$ and $G_{2}$. The join $G_{1}\vee G_{2}$ is the graph obtained from $G_{1}+G_{2}$ by adding all possible edges between $V(G_{1})$ and $V(G_{2})$. For other concepts, please refer to [3].
For a connected graph $G$ of order $n$, the distance between vertices $v_{i}$ and $v_{j}$ denoted by $d_{i,j}$, is the length of a shortest path from $v_{i}$ to $v_{j}$ in $G$. The distance matrix $D(G)=(d_{ij})$ of $G$ is a square matrix of order $n$, where $d_{ij}$ is the distance from vertex $v_{i}$ to vertex $v_{j}$. The largest eigenvalue of $D(G)$, denoted by $\mu(G)$, is called the distance spectral radius of $G$.
A fractional matching of a connected graph $G$ is a function $f$ giving each edge a number in $[0,1]$ such that for each $v_{i}\in V(G)$, $\sum_{e\in E_{G}(v_{i})}f(e)\leq 1$, where $E_{G}(v_{i})=\{e:e\in E(G)\ \textrm{and}\ e\ \textrm{ is incident with}\ v_{i}\}$. Let $\alpha_{f}(G)$ denote the fractional matching number of $G$ which is defined as $\alpha_{f}(G)=\max\{\sum_{e\in E(G)}f(e):f\ \textrm{is a fractional matching of}\ G\}$. Summing the inequality constraints for all vertices has $2\sum_{e\in E(G)}f(e)\leq n$, thus $\alpha_{f}(G)\leq\frac{n}{2}$. Therefore, a fractional matching of $G$ is called a fractional perfect matching if $\sum_{e\in E(G)}f(e)=\frac{n}{2}$.
Recently, the relationship between the fractional matching number and the eigenvalues of graphs has been investigated by several researchers. For example, O [7] determined the connections between the spectral radius of an $n$-vertex connected graph with minimum degree and its fractional matching number, and gave a lower bound on the fractional matching number in terms of the spectral radius and minimum degree. Xue et al. [11] discussed the relations between the fractional matching number and the Laplacian spectral radius of a graph, and obtained some lower bounds on the fractional matching number. Pan et al. [8] studied the existence of fractional perfect matchings for graphs with given order and minimum degree in terms of the signless Laplacian spectral radius. Yan et al. [12] gave some lower bounds of distance Laplacian spectral radii of $n$-vertex graphs in terms of fractional matching number. Lou et al. [5] obtained a tight lower bound of the spectral radius to guarantee the fractional matching number more than $\frac{n-k}{2}$ in a graph with minimum degree $\delta$. Motivated by [5], in this paper we firstly investigate the relations between the distance spectral radius of a graph and its fractional matching number.
For a set $\{G_{1},G_{2},G_{3},\cdots\}$ of graphs, an $\{G_{1},G_{2},G_{3},\cdots\}$-factor of a graph $G$ is a spanning subgraph of $G$ each component of which is isomorphic to one of $\{G_{1},G_{2},G_{3},\cdots\}$. Hence, if $H$ is an $\{G_{1},G_{2},G_{3},\cdots\}$-factor of graph $G$, then $H$ is a subgraph of $G$ such that $V(H)=V(G)$ and each component of $H$ is contained in $\{G_{1},G_{2},G_{3},\cdots\}$. Specifically, the $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor is also called a star factor of $G$.
The factor of graphs have received a lot of attention of researchers and its theories are well developed in recent years. In [10], Tutte presented a sufficient and necessary condition for the existence of a $\{K_{2},{C_{k}}\}$-factor. In [1], Amahashi and Kano gave a necessary and sufficient condition of the existence of a $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor in a graph, where $k\geq 2$ is an integer. In [6], Miao and Li established a lower bound on the size of $G$ to guarantee that $G$ contains a star factor. They also determined an upper bound on the spectral radius of $G$ to ensure that $G$ has a star factor. In [5], Lou et al. provided a tight spectral radius condition for the existence of a $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor with $k\geq 2$ in a graph with minimum degree $\delta$. Motivated by [5], in this paper we also extend their results to the distance spectral radius.
The rest of this paper is divided into the following sections. In Section 2, we give some concepts and lemmas to prove the theorems in the following sections. In Section 3, we establish a sharp upper bound for the distance spectral radius to guarantee $\alpha_{f}(G)>\frac{n-k}{2}$ in graph $G$ with given minimum degree $\delta$. In Section 4, we give a sharp upper bound on distance spectral radius of graph $G$ with given minimum degree $\delta$ to ensure that $G$ has a $\{K_{2},\{C_{k}\}\}$-factor, where $k\geq 3$ is an integer. In addition, we also obtain a sharp upper bounds on distance spectral radius for the existence of a $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor with $k\geq 2$ in a graph $G$ with given minimum degree $\delta$.
2 Preliminaries
In this section, we give some concepts and lemmas which will be used in the follows. Firstly, we give some lemmas about fractional matching number and the sufficient and necessary condition for the existence of $\{K_{2},\{C_{k}\}\}$-factor and $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor in a graph. For $S\subseteq V(G)$, we use $i(G-S)$ denotes the number of isolated vertices in the graph $G-S$.
Lemma 2.1.
([9])  Let $G$ be a graph of order $n$. Then
$\alpha_{f}(G)=\frac{1}{2}(n-\max\{i(G-S)-|S|:\textrm{for all}\ S\subseteq V(G)\}).$
Lemma 2.2.
([10])  Let $G$ be a graph of order $n$ and $k\geq 3$ be an integer. Then $G$ has a $\{K_{2},\{C_{k}\}\}$-factor if and only if $i(G-S)\leq|S|$ for every $S\subseteq V(G)$.
Lemma 2.3.
([1])  Let $G$ be a graph of order $n$ and $k\geq 2$ be an integer. Then $G$ has a $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor if and only if $i(G-S)\leq k|S|$ for every $S\subseteq V(G)$.
Nonnegative matrix theory plays an important role in the study of spectral graph theory. Next we will introduce some concepts and theories of nonnegative matrices.
Definition 2.4.
([2]) Let $A=(a_{ij})$ and $B=(b_{ij})$ be two $n\times n$ matrices. If $a_{ij}\leq b_{ij}$ for all $i$ and $j$, then $A\leq B$. If $A\leq B$ and $A\neq B$, then $A<B$. If $a_{ij}<b_{ij}$ for all $i$ and $j$, then $A\ll B$.
Lemma 2.5.
([2])  Let $A$ and $B$ be two $n\times n$ nonnegative matrices with the spectral radius $\rho(A)$ and $\rho(B)$, respectively. If $A\leq B$, then $\rho(A)\leq\rho(B)$. In addition, if $A<B$ and $B$ is irreducible, then $\rho(A)<\rho(B)$.
Lemma 2.6.
([2])  Let $m<n$, $A$ and $B$ be $n\times n$ and $m\times m$ nonnegative matrices with the spectral radius $\rho(A)$ and $\rho(B)$, respectively. If $B$ is a principal submatrix of $A$, then $\rho(B)\leq\rho(A)$. In addition, if $A$ is irreducible, then $\rho(B)<\rho(A)$.
By Lemma 2.5, we can get an important result about the distance radius of graph $G$.
Lemma 2.7.
([4])  Let $e$ be an edge of $G$ such that $G-e$ is connected. Then $\mu(D(G))<\mu(D(G-e))$.
Finally, we give the definition of the equitable quotient matrix and its some properties.
Definition 2.8.
([13]) Let $M$ be a complex matrix of order $n$ described in the following block form
$M=\begin{bmatrix}M_{11}&\cdots&M_{1t}\\ \vdots&\ddots&\vdots\\ M_{t1}&\cdots&M_{tt}\end{bmatrix},$
where the blocks $M_{ij}$ are $n_{i}\times n_{j}$ matrices for any $1\leq i,j\leq t$ and $n=n_{1}+n_{2}+\cdots+n_{t}.$ For $1\leq i,j\leq t$, let $b_{ij}$ denote the average row sum of $M_{ij}$, i.e. $b_{ij}$ is the sum of all entries in $M_{ij}$ divided by the number of rows. Then $B(M)=(b_{ij})$(or simply $B$) is called the quotient matrix of $M$. If, in addition, for each pair $i$, $j$, $M_{ij}$ has a constant row sum, then $B$ is called the equitable quotient matrix of $M$.
Lemma 2.9.
([13]) Let $B$ be the equitable quotient matrix of $M$. In addition, let $M$ be a nonnegative matrix. Then the spectral radius relation satisfies $\rho(B)=\rho(M)$.
3 A sharp upper bound for the distance spectral radius to guarantee $\alpha_{f}(G)>\frac{n-k}{2}$ with given minimum degree $\delta$.
First of all, we give a theorem which was proposed by Scheinerman and Ullman [9]. This theorem plays an important role in the fractional matching of graphs.
Theorem 3.1.
([9])  Let $G$ be a graph of order $n$.
(a) Any fractional matching satisfies $\alpha_{f}(G)\leq\frac{n}{2}$.
(b) $2\alpha_{f}(G)$ is an integer.
Since $\alpha_{f}(G)\leq\frac{n}{2}$, it is natural to consider that can we find a condition for distance spectral radius that makes the fractional matching number of a graph $G$ with given minimum degree more than $\frac{n-k}{2}$? In addition, can we characterize the corresponding spectral extremal graphs? Based on the above considerations, we give the following theorem.
Theorem 3.2.
Let $G$ be a graph of order $n>9k+10\delta+2$ with minimum degree $\delta$, where $0<k<n$ is an integer. If $\mu(G)\leq\mu(K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1}))$, then $\alpha_{f}(G)>\frac{n-k}{2}$ unless $G\cong K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1}).$
Proof.
Let $G$ be a graph of order $n>9k+10\delta+2$. The minimum degree of $G$ denoted by $\delta$. By way of contradiction assume that the fractional matching number $\alpha_{f}(G)\leq\frac{n-k}{2}$. Then by Lemma 2.1, there exists a vertex subset $S\subseteq V(G)$ such that $\alpha_{f}(G)=\frac{1}{2}(n-\max\{i(G-S)-|S|:\forall S\subseteq V(G)\})$. Thus we have $i(G-S)-|S|\geq k$. The set of isolated vertices in $G-S$ denoted by $T$. Let $|S|=s$ and $|T|=t$. Then we have $t=i(G-S)\geq s+k$. Since $s+t\leq n$, we have $s\leq\frac{n-k}{2}$. It is easy to find that $N_{G}(T)\subseteq S$, thus we have $s\geq\delta$. Let $G_{1}=K_{s}\vee(K_{n-2s-k}+(s+k)K_{1})$. Obviously, $G$ is a spanning subgraph of $G_{1}$. By Lemma 2.7, we have
$\mu(G)\geq\mu(G_{1}),$ (1)
where equality holds if and only if $G\cong G_{1}$. Since $\delta\leq s\leq\frac{n-k}{2}$, we will discuss the proof in two ways according to the value of $s$. Case 1: $s=\delta.$ Then $G_{1}=K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1})$. By (1), we have
$\mu(G)\geq\mu(G_{1})=\mu(K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1})).$
What’s more, according to the assumed condition $\mu(G)\leq\mu(K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1}))$ and $G$ is a spanning subgraph of $K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1})$, we conclude that $G\cong K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1})$. Obviously, deleting vertices in $K_{n-2\delta-k}$ and $(\delta+k)K_{1}$ does not cause graph $G$ to become disconnected. Hence,
$\max\{i(G-S)-|S|:\forall S\subseteq V(G)\}=i(G-V(K_{\delta}))-|V(K_{\delta})|=k.$
Thus, with the help of Lemma 2.1, through a simple calculation, we have $\alpha_{f}(G)=\frac{n-k}{2}$. Hence, if $s=\delta$, we have $G\cong K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1})$. Case 2: $\delta<s\leq\frac{n-k}{2}.$ Let $G_{1}=K_{s}\vee(K_{n-2s-k}+(s+k)K_{1})$. We divide $V(G_{1})$ into three parts: $V(K_{s})$, $V(K_{n-2s-k})$, $V((s+k)K_{1})$. Then the distance matrix of $G_{1}$, denoted by $D(G_{1})$, is
$\begin{bmatrix}(J-I)_{s\times s}&J_{s\times(n-2s-k)}&J_{s\times(s+k)}\\ J_{(n-2s-k)\times s}&(J-I)_{(n-2s-k)\times(n-2s-k)}&2J_{(n-2s-k)\times(s+k)}\\ J_{(s+k)\times s}&{2J}_{(s+k)\times(n-2s-k)}&2(J-I)_{(s+k)\times(s+k)}\end{bmatrix},$
where $J$ denotes the all-one matrix and $I$ denotes the identity square matrix. Then we use $M_{s}$ denote the equitable quotient matrix of the distance matrix $D(G_{1})$ for the partition $V(K_{s})\cup V(K_{n-2s-k})\cup V((s+k)K_{1})$. Thus
$M_{s}=\begin{bmatrix}s-1&n-2s-k&s+k\\ s&n-2s-k-1&2(s+k)\\ s&2(n-2s-k)&2(s+k-1)\end{bmatrix}.$
Through a simple calculation, we get the characteristic polynomial of $M_{s}$ is
$f_{s}(x)$ $=x^{3}+(-s-n-k+4)x^{2}+(5s^{2}-2ns+7ks-s-2kn-3n+2k^{2}-k+5)x$
$\ \ \ -2s^{3}+(n-3k+5)s^{2}+(kn-2n-k^{2}+7k)s-2kn-2n+2k^{2}+2.$
Let $\lambda_{1}(M_{s})$ denote the largest real root of the equation $f_{s}(x)=0$. By Lemma 2.9, we have that $\mu(G_{1})=\lambda_{1}(M_{s})$. Obviously, by replacing $s$ with $\delta$, we can get a equitable quotient matrix $M_{\delta}$ of $G_{2}=K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1})$. Similarly, we can get the characteristic polynomial $f_{\delta}(x)$ of $M_{\delta}$ and $\mu(G_{2})=\lambda_{1}(M_{\delta})$ is the largest real root of the equation $f_{\delta}(x)=0$. Then we have
$f_{s}(x)-f_{\delta}(x)$ $=(\delta-s)[x^{2}+(2n+1-5(\delta+s)-7k)x+(2s-n+2\delta+3k-5)s$
$\ \ \ \ +(2\delta-n+3k-5)\delta+2n-kn+k^{2}-7k].$
Since $G_{1}$ and $G_{2}$ are spanning subgraphs of $K_{n}$, by lemma 2.7, we have $\mu(G_{1})>\mu(K_{n})=n-1$ and $\mu(G_{2})>\mu(K_{n})=n-1$. Then we will proof that $f_{s}(x)-f_{\delta}(x)<0$ for $x\in[n-1,+\infty)$. Since $\delta<s$, we only need to proof $g(x)>0$, where
$g(x)=x^{2}+(2n+1-5(\delta+s)-7k)x+(2s-n+2\delta+3k-5)s+(2\delta-n+3k-5)\delta+2n-kn+k^{2}-7k.$
Since $s\leq\frac{n-k}{2}$, the symmetry axis of $g(x)$ is
$\tilde{x}$ $=\frac{5(\delta+s)+7k-2n-1}{2}$
$=\frac{5}{2}s+\frac{5}{2}\delta+\frac{7}{2}k-n-\frac{1}{2}$
$\leq\frac{5}{4}(n-k)+\frac{5}{2}\delta+\frac{7}{2}k-n-\frac{1}{2}$
$=\frac{1}{4}n+\frac{5}{2}\delta+\frac{9}{4}k-\frac{1}{2}.$
Since $n>9k+10\delta+2>3k+\frac{10}{3}\delta+\frac{2}{3}$ and $n>3k+\frac{10}{3}\delta+\frac{2}{3}\iff\frac{1}{4}n+\frac{5}{2}\delta+\frac{9}{4}k-\frac{1}{2}<n-1$, we have $\frac{5(\delta+s)+7k-2n-1}{2}<n-1$, which implies $g(x)$ is increasing with respect to $x\in[n-1,+\infty)$. Then
$g(x)\geq g(n-1)=2s^{2}+(3k+2\delta-6n)s+3n^{2}+2\delta^{2}+(3k-6n)\delta-n-8kn+k^{2}.$
Let
$h(s)=g(n-1)=2s^{2}+(3k+2\delta-6n)s+3n^{2}+2\delta^{2}+(3k-6n)\delta-n-8kn+k^{2}.$
Note that $\delta<s\leq\frac{n-k}{2}$, we have
$\frac{dh}{ds}$ $=4s+3k+2\delta-6n$
$\leq 2n-2k+3k+2\delta-6n$
$=-4n+k+2\delta<0.$
Hence $h(s)$ is decreasing with respect to $s\in[\delta+1,\frac{n-k}{2}]$. By a simple calculation, we have
$h(s)\geq h(\frac{n-k}{2})$ $=\frac{1}{2}[n^{2}-(10\delta+9k+2)n+4\delta^{2}+4k\delta]$
$>\frac{1}{2}[n^{2}-(10\delta+9k+2)n].$
Since $n>10\delta+9k+2$, we have $h(s)>0$. Hence, we have $g(x)\geq g(n-1)=h(s)>0$, which implies $f_{s}(x)<f_{\delta}(x)$ for $x\in[n-1,+\infty)$. Since $\min\{\mu(G_{1}),\mu(G_{2})\}>n-1$, we have $\mu(G_{1})>\mu(G_{2})$. Thus
$\mu(G)\geq\mu(G_{1})>\mu(G_{2})=\mu(K_{\delta}\vee(K_{n-2\delta-k}+(\delta+k)K_{1}),$
a contradiction. This completes the proof. ∎
Let $k=1$, we can get a distance spectral radius condition in a graph $G$ of order $n$ with minimum degree $\delta$ such that $\alpha_{f}(G)>\frac{n-1}{2}$. However, by Theorem 3.1, we have $\alpha_{f}(G)\leq\frac{n}{2}$ and $2\alpha_{f}(G)$ is an integer. Thus, we can obtain the following corollary about fractional perfect matching based on the distance spectral radius.
Corollary 3.3.
Let $G$ be a graph of order $n>11+10\delta$ with minimum degree $\delta$. If $\mu(G)\leq\mu(K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1}))$, then $G$ contains a fractional perfect matching unless $G\cong K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1}).$
4 A sharp upper bound on distance spectral radius for the existence of a $\{K_{2},\{C_{k}\}\}$-factor with $k\geq 3$ and $\{K_{1,1},K_{1,2},\cdots,$ $K_{1,k}\}$-factor with $k\geq 2$ in a graph with given minimum degree $\delta$
Tutte [10] gave a sufficient and necessary condition for a graph contain a $\{K_{2},\{C_{k}\}\}$-factor with $k\geq 3$. In this section, we firstly obtain a distance spectral radius condition that makes a graph $G$ with given minimum degree $\delta$ has a $\{K_{2},\{C_{k}\}\}$-factor.
Theorem 4.1.
Let $G$ be a graph of order $n>11+10\delta$ with minimum degree $\delta$, where $k\geq 3$ is an integer. If $\mu(G)\leq\mu(K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1}))$, then $G$ has a $\{K_{2},\{C_{k}\}\}$-factor unless $G\cong K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1}).$
Proof.
Let $G$ be a graph of order $n>11+10\delta$. The minimum degree of $G$ denoted by $\delta$. By way of contradiction assume that $G$ has no a $\{K_{2},\{C_{k}\}\}$-factor. Then by Lemma 2.2, there exists a vertex subset $S\subseteq V(G)$ such that $i(G-S)-|S|\geq 1$. The set of isolated vertices in $G-S$ denoted by $T$. Let $|S|=s$ and $|T|=t$. Then we have $t=i(G-S)\geq s+1$. Since $s+t\leq n$, we have $s\leq\frac{n-1}{2}$. It is easy to find that $N_{G}(T)\subseteq S$, thus we have $s\geq\delta$. Let $G_{1}=K_{s}\vee(K_{n-2s-1}+(s+1)K_{1})$. Obviously, $G$ is a spanning subgraph of $G_{1}$. By Lemma 2.7, we have
$\mu(G)\geq\mu(G_{1}),$ (2)
where equality holds if and only if $G\cong G_{1}$. Since $\delta\leq s\leq\frac{n-1}{2}$, we will discuss the proof in two ways according to the value of $s$.
Case 1: $s=\delta$.
Then $G_{1}=K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1})$. By (2), we have
$\mu(G)\geq\mu(G_{1})=\mu(K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1})).$
Moreover, according to the assumed condition $\mu(G)\leq\mu(K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1}))$ and $G$ is a spanning subgraph of $K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1})$, we conclude that $G\cong K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1})$. Note that the vertices of $(\delta+1)K_{1}$ are only adjacent to $\delta$ vertices of $K_{\delta}$. Let $G_{0}$ be a connected spanning subgraph of $K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1})$, it is easy to find that $G_{0}$ is neither isomorphic to $K_{2}$ nor to $C_{k}$. For any disconnected spanning subgraph of $K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1})$, we can always find a component isomorphic to a single vertex set $K_{1}$. Hence, $K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1})$ does not have a $\{K_{2},\{C_{k}\}\}$-factor.
Case 2: $\delta<s\leq\frac{n-1}{2}$.
Let $G_{1}=K_{s}\vee(K_{n-2s-1}+(s+1)K_{1}$. We divide $V(G_{1})$ into three parts: $V(K_{s})$, $V(K_{n-2s-1})$, $V((s+1)K_{1})$. Then we use $M_{s}$ denotes the equitable quotient matrix of the distance matrix $D(G_{1})$ for the partition $V(K_{s})\cup V(K_{n-2s-1})\cup V((s+1)K_{1})$ of $V(G_{1})$. Thus
$M_{s}=\begin{bmatrix}s-1&n-2s-1&s+1\\ s&n-2s-2&2(s+1)\\ s&2(n-2s-1)&2s\end{bmatrix}.$
By a simple calculation, we get the characteristic polynomial of $M_{s}$ is
$f_{s}(x)=x^{3}+(3-s-n)x^{2}+(5s^{2}-2ns+6s-5n+6)x-2s^{3}+(n+2)s^{2}+(6-n)s-4n+4.$
Let $\lambda_{1}(M_{s})$ denote the largest real root of the equation $f_{s}(x)=0$. By Lemma 2.9, we have that $\mu(G_{1})=\lambda_{1}(M_{s})$. Obviously, by replacing $s$ with $\delta$, we can get a equitable quotient matrix $M_{\delta}$ of $G_{2}=K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1})$. Similarly, we can get the characteristic polynomial $f_{\delta}(x)$ of $M_{\delta}$ and $\mu(G_{2})=\lambda_{1}(M_{\delta})$ is the largest real root of the equation $f_{\delta}(x)=0$. Then we have
$f_{s}(x)-f_{\delta}(x)=(\delta-s)[x^{2}+(2n-5(\delta+s)-6)x+(2s-n+2\delta-2)s+(2\delta-n-2)\delta+n-6].$
Since $G_{1}$ and $G_{2}$ are spanning subgraphs of $K_{n}$, by lemma 2.7, we have $\mu(G_{1})>\mu(K_{n})=n-1$ and $\mu(G_{2})>\mu(K_{n})=n-1$. Then we will proof that $f_{s}(x)-f_{\delta}(x)<0$ for $x\in[n-1,+\infty)$. Since $\delta<s$, we only need to proof $g(x)>0$, where
$g(x)=x^{2}+(2n-5(\delta+s)-6)x+(2s-n+2\delta-2)s+(2\delta-n-2)\delta+n-6.$
Since $s\leq\frac{n-1}{2}$, the symmetry axis of $g(x)$ is
$\tilde{x}$ $=\frac{5(\delta+s)+6-2n}{2}$
$=\frac{5}{2}s+\frac{5}{2}\delta+\frac{7}{2}-n-\frac{1}{2}$
$\leq\frac{5}{4}(n-1)+\frac{5}{2}\delta+\frac{7}{2}-n-\frac{1}{2}$
$=\frac{1}{4}n+\frac{5}{2}\delta+\frac{7}{4}.$
Since $n>10\delta+11>\frac{10}{3}\delta+\frac{11}{3}$ and $n>\frac{10}{3}\delta+\frac{11}{3}\iff\frac{1}{4}n+\frac{5}{2}\delta+\frac{7}{4}<n-1$, we have $\frac{5(\delta+s)+6-2n}{2}<n-1$, which implies $g(x)$ is increasing with respect to $x\in[n-1,+\infty)$. Then
$g(x)\geq g(n-1)=2s^{2}+(3+2\delta-6n)s+3n^{2}+2\delta^{2}+(3-6n)\delta-9n+1.$
Let
$h(s)=g(n-1)=2s^{2}+(3+2\delta-6n)s+3n^{2}+2\delta^{2}+(3-6n)\delta-9n+1.$
Note that $\delta<s\leq\frac{n-1}{2}$, we have
$\frac{dh}{ds}$ $=4s+3+2\delta-6n$
$\leq 2n-2+3+2\delta-6n$
$=-4n+1+2\delta<0.$
Hence $h(s)$ is decreasing with respect to $s\in[\delta+1,\frac{n-1}{2}]$. By a simple calculation, we have
$h(s)\geq h(\frac{n-1}{2})$ $=\frac{1}{2}[n^{2}-(10\delta+11)n+4\delta^{2}+4\delta]$
$>\frac{1}{2}[n^{2}-(10\delta+11)n].$
Since $n>10\delta+11$, we have $h(s)>0$. Hence, we have $g(x)\geq g(n-1)=h(s)>0$, which implies $f_{s}(x)<f_{\delta}(x)$ for $x\in[n-1,+\infty)$. Since $\min\{\mu(G_{1}),\mu(G_{2})\}>n-1$, we have $\mu(G_{1})>\mu(G_{2})$. Thus
$\mu(G)\geq\mu(G_{1})>\mu(G_{2})=\mu(K_{\delta}\vee(K_{n-2\delta-1}+(\delta+1)K_{1}),$
a contradiction. This completes the proof. ∎
In the following, we will present a distance spectral radius condition to guarantee the existence of $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor in a graph $G$ with given minimum degree $\delta$.
Theorem 4.2.
Let $G$ be a graph of order $n>\frac{3+5k}{k^{2}}+3+(\frac{3}{k}+5+2k)\delta$ with minimum degree $\delta$, where $k\geq 2$ is an integer. If $\mu(G)\leq\mu(K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1}))$, then $G$ has a $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor unless $G\cong K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1}).$
Proof.
Let $G$ be a graph of order $n>\frac{3+5k}{k^{2}}+3+(\frac{3}{k}+5+2k)\delta$. The minimum degree of $G$ denoted by $\delta$. By way of contradiction assume that $G$ has no a $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor. Then by Lemma 2.3, there exists a vertex subset $S\subseteq V(G)$ such that $i(G-S)-k|S|\geq 1$. The set of isolated vertices in $G-S$ denoted by $T$. Let $|S|=s$ and $|T|=t$. Then we have $t=i(G-S)\geq ks+1$. Since $s+t\leq n$, we have $s\leq\frac{n-1}{k+1}$. It is easy to find that $N_{G}(T)\subseteq S$, thus we have $s\geq\delta$. Let $G_{1}=K_{s}\vee(K_{n-ks-s-1}+(ks+1)K_{1})$. Obviously, $G$ is a spanning subgraph of $G_{1}$. By Lemma 2.7, we have
$\mu(G)\geq\mu(G_{1}),$ (3)
where equality holds if and only if $G\cong G_{1}$. Since $\delta\leq s\leq\frac{n-1}{k+1}$, we will discuss the proof in two ways according to the value of $s$. Case 1: $s=\delta.$ Then $G_{1}=K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1})$. By (3), we have
$\mu(G)\geq\mu(G_{1})=\mu(K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1})).$
Furthermore, according to the assumed condition $\mu(G)\leq\mu(K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1}))$ and $G$ is a spanning subgraph of $K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1})$, we conclude that $G\cong K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1})$. Note that the vertices of $(k\delta+1)K_{1}$ are only adjacent to $\delta$ vertices of $K_{\delta}$. Let $G_{0}$ be a connected spanning subgraph of $K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1})$, it is easy to find that $G_{0}$ is not isomorphic to any graph in $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$. For any disconnected spanning subgraph of $K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1})$, we can always find a component isomorphic to a single vertex set $K_{1}$. Hence $K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1})$ does not contain a spanning subgraph each component of which is contained in $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$. This implies that $K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1})$ does not have a $\{K_{1,1},K_{1,2},\cdots,K_{1,k}\}$-factor. Case 2: $\delta<s\leq\frac{n-1}{k+1}.$ Let $G_{1}=K_{s}\vee(K_{n-ks-s-1}+(ks+1)K_{1})$. We divide $V(G_{1})$ into three parts: $V(K_{s})$, $V(K_{n-ks-s-1})$, $V((ks+1)K_{1})$. Then we use $M_{s}$ for the equitable quotient matrix of the distance matrix $D(G_{1})$ for the partition $V(K_{s})\cup V(K_{n-ks-s-1})\cup V((ks+1)K_{1})$ of $V(G_{1})$. Thus
$M_{s}=\begin{bmatrix}s-1&n-ks-s-1&ks+1\\ s&n-ks-s-2&2(ks+1)\\ s&2(n-ks-s-1)&2ks\end{bmatrix}.$
By a simple calculation, we get the characteristic polynomial of $M_{s}$ is
$f_{s}(x)$ $=x^{3}+(-ks-n+3)x^{2}+(2k^{2}s^{2}+3ks^{2}-2kns+3ks+3s-5n+6)x-(k^{2}+k)s^{3}$
$\ \ \ \ +(kn+2k^{2}+k-1)s^{2}+(n-2kn+4k+2)s+4-4n.$
Let $\lambda_{1}(M_{s})$ denote the largest real root of the equation $f_{s}(x)=0$. By Lemma 2.9, we have that $\mu(G_{1})=\lambda_{1}(M_{s})$. Obviously, by replacing $s$ with $\delta$, we can get a equitable quotient matrix $M_{\delta}$ of $G_{2}=K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1})$. Similarly, we can get the characteristic polynomial $f_{\delta}(x)$ of $M_{\delta}$ and $\mu(G_{2})=\lambda_{1}(M_{\delta})$ is the largest real root of the equation $f_{\delta}(x)=0$. Then we have
$f_{s}(x)-f_{\delta}(x)$ $=(\delta-s)[kx^{2}+(-2k^{2}s-3ks+2kn-2k^{2}\delta-3k\delta-3k-3)x$
$\ \ \ +(k^{2}s+ks-kn+k^{2}\delta+k\delta-2k^{2}-k+1)s$
$\ \ \ +(-kn+k^{2}\delta+k\delta-2k^{2}-k+1)\delta-n+2kn-4k-2].$
Since $G_{1}$ and $G_{2}$ are spanning subgraphs of $K_{n}$, by lemma 2.7, we have $\mu(G_{1})>\mu(K_{n})=n-1$ and $\mu(G_{2})>\mu(K_{n})=n-1$. Then we will proof that $f_{s}(x)-f_{\delta}(x)<0$ for $x\in[n-1,+\infty)$. Since $\delta<s$, we only need to proof $g(x)>0$, where
$g(x)=kx^{2}+(-2k^{2}s-3ks+2kn-2k^{2}\delta-3k\delta-3k-3)x+(k^{2}s+ks-kn+k^{2}\delta+k\delta-2k^{2}-k+1)s.$
Since $s\leq\frac{n-1}{k+1}$ and $k\geq 2$, the symmetry axis of $g(x)$ is
$\tilde{x}$ $=-\frac{(-2k^{2}s-3ks+2kn-2k^{2}\delta-3k\delta-3k-3)}{2k}$
$=\frac{3}{2k}+\frac{3}{2}+(\frac{3}{2}+k)\delta+\frac{3}{2}s+ks-n$
$\leq\frac{9}{4}+(\frac{3}{2}+k)\delta+(\frac{3}{2}+k)s-n$
$\leq\frac{9}{4}+(\frac{3}{2}+k)\delta+(\frac{3}{2}+k)\frac{n-1}{k+1}-n$
$=\frac{5}{4}+(\frac{3}{2}+k)\delta+\frac{1}{2(k+1)}n-\frac{1}{2(k+1)}.$
Since $n>\frac{3+5k}{k^{2}}+3+(\frac{3}{k}+5+2k)\delta>\frac{1}{2k+1}[\frac{9}{2}(k+1)-1]+(\frac{3}{2}+k)\frac{2(k+1)}{2k+1}\delta$ and $n>\frac{1}{2k+1}[\frac{9}{2}(k+1)-1]+(\frac{3}{2}+k)\frac{2(k+1)}{2k+1}\delta\iff\frac{5}{4}+(\frac{3}{2}+k)\delta+\frac{1}{2(k+1)}n-\frac{1}{2(k+1)}<n-1$, we have $-\frac{(-2k^{2}s-3ks+2kn-2k^{2}\delta-3k\delta-3k-3)}{2k}<n-1$, which implies $g(x)$ is increasing with respect to $x\in[n-1,+\infty)$. Then
$g(x)\geq g(n-1)$ $=(k^{2}+k)s^{2}+(1+2k+k\delta+k^{2}\delta-4kn-2k^{2}n)s+3kn^{2}$
$\ \ \ +k(k+1)\delta^{2}+(1+2k-4kn-2k^{2}n)\delta-4n-5kn+1.$
Let
$h(s)=g(n-1)$ $=(k^{2}+k)s^{2}+(1+2k+k\delta+k^{2}\delta-4kn-2k^{2}n)s+3kn^{2}$
$\ \ \ +k(k+1)\delta^{2}+(1+2k-4kn-2k^{2}n)\delta-4n-5kn+1.$
Note that $\delta<s\leq\frac{n-1}{k+1}$ and $n>\frac{3+5k}{k^{2}}+3+(\frac{3}{k}+5+2k)\delta$, we have
$\frac{dh}{ds}$ $=(2k^{2}+2k)s+1+2k+k\delta+k^{2}\delta-4kn-2k^{2}n$
$\leq 2k(n-1)+1+2k+k\delta+k^{2}\delta-4kn-2k^{2}n$
$=-2kn+1+k(k+1)\delta-2k^{2}n<0.$
Hence $h(s)$ is decreasing with respect to $s\in[\delta+1,\frac{n-1}{k+1}]$. By a simple calculation, we have
$h(s)\geq h(\frac{n-1}{k+1})$ $=\frac{1}{k+1}[k^{2}n^{2}-(3+5k+3k^{2}+(3k+5k^{2}+2k^{3})\delta)n$
$\ \ \ +k^{3}\delta^{2}+2\delta^{2}k^{2}+k\delta^{2}+k^{2}\delta+2k\delta+1]$
$\ \ \ >\frac{1}{k+1}[k^{2}n^{2}-(3+5k+3k^{2}+(3k+5k^{2}+2k^{3})\delta)n].$
Since $n>\frac{3+5k}{k^{2}}+3+(\frac{3}{k}+5+2k)\delta$, we have $h(s)>0$. Hence, we have $g(x)\geq g(n-1)=h(s)>0$, which implies $f_{s}(x)<f_{\delta}(x)$ for $x\in[n-1,+\infty)$. Since $\min\{\mu(G_{1}),\mu(G_{2})\}>n-1$, we have $\mu(G_{1})>\mu(G_{2})$. Thus
$\mu(G)\geq\mu(G_{1})>\mu(G_{2})=\mu(K_{\delta}\vee(K_{n-k\delta-\delta-1}+(k\delta+1)K_{1}),$
a contradiction. This completes the proof. ∎